Sampling plan

• Great to be trying a phone survey, and great that you found phones.whitepages.com/412.

• Don’t call cell #’s. Generally speaking people have to call for incoming minutes; this is a burden you shouldn’t impose.

• You might want to stratify by major region of the city (Downtown, Oakland/campus area, east end, South side, etc.) but I advise against stratifying by neighborhood or telephone exchange.

– Actually, I \underline{strongly suggest} you do a srs, using a method for phone numbers similar to what i suggested for cbook: for each xxx-yyyy number randomly choose an (xxx,yyyy) pair where xxx is a random choice of one of the non-cellphone exchanges in 412, and yyyy is a random # from 0000 to 9999.

– since you are asking for neighborhood in your survey anyway you can get some regional information without stratifying, and it will greatly simplify your data collection.

– you will need to keep track of not-answered phones, etc, in order to followup nonresponses, compute a nonresponse rate, etc.

PLEASE PROVIDE AN ESTIMATED SAMPLE SIZE FOR YOUR SURVEY. THAT IS PART OF THIS ASSIGNMENT!

Questionnaire and interview script

• Your questionnaire needs to be reformatted into an interview script, so that you make the same introduction, ask for the same type of person, and ask the same questions in the same way, with every household you successfully contact.

you need to set the script up as soon as possible and try it out, both for “successful” interviews and for ones that go off-track, so you can anticipate some of the possible problems that will arise, and so that you can refine and clarify your questions.

• I am attaching a couple of documents that have advice on how to set up the beginning of a phone interview script, and what some of the issues are.

Generally a script should have the following elements (some example text is provided, it will need to be improved for your script!)

Introduction Hello, I am XXX, a student at Carnegie Mellon University, calling you as part of a survey for our class ZZZZ.

– Is this a home, or apartment? [if not, apologize, thank the person and end call ]

– I would like to speak to XXXX (you have to choose who you want to talk to, and make the same choice for every household. if that kind of person is not avail, apologize, thanks, and end call)

when person you want gets on the phone:

– Hello, I am XXX, a student at Carnegie Mellon University, calling you as part of a survey for our class ZZZZ.

(if asked at any time during survey) Please feel free to contact our instructor, Brian Junker, 412 268 8874, at Carnegie Mellon University, with any questions you may have about this survey or our class.

– your phone # was randomly selected for this survey. The survey is about bicycle usage in pgh. The survey has XX questions and will take approximately YY minutes to complete.

– Your phone number will not be associated with your responses, and the information from the survey will only be reported in summarized form, and your responses will be kept confidential. Your participation is completely voluntary and you may refuse to answer any question you wish, and you may stop at any time.

– Do you wish to start the survey? [if no, sorry, thanks, end call]

Body of survey This is where you ask the survey questions.

– Every question should be carefully worded, and practiced by you, for clear speaking, simple and easy to follow formal English, etc.

– Structure the survey so that similar questions are grouped together. If possible a “key question” can eliminate others.

∗ For example, “do you own or use a bike”… “No” eliminates most of the rest

of the survey!

– Anticipate problems and have something in the script to address them.

∗ For example: does the person not understand something (what is a neighbor-

hood? what is a sector of the economy?)? Give examples.

– Have a plan for recording the answers as you get them. Eg.

∗ design a paper form that you can xerox copies of for each call. it should be easy to read off the script from the form, and it should be easy to write answers — even unusual ones! — in the space after each question.

∗ if you prefer do the same thing on a laptop with an appropriately formatted msword or excel file – make a copy of the file for each phone contact, and fill it out as you do the interview. Remember to save the file!

The person doing the interview should practice practice practice reading the questions and recording the data so that it goes smoothly and easily when you are talking to a “real” respondent.

∗ Always polite

∗ Always clear

∗ Easy to understand speaking

∗ Get all the data

∗ don’t ask the interviewee to repeat unless you really have to

∗ unfazed by anything that might come up (e.g being cussed at)

∗ Always polite

∗ Always clear

∗ Easy to understand speaking

End of survey – whether respondent completed all the questions or not! Thank you very much for your time. If you would like to find out the results of this survey [tell them how, e.g. maybe download your report from a website after date XXX?]

Please feel free to contact our instructor, Brian Junker, 412 268 8874, at Carnegie Mellon University, with any questions you may have about this survey or our class.